OpenStudy (anonymous):

The radius of a hydrogen atom is about 0.37Å, and the average radius of the earth's orbit around the sun is about 1.5x10^8 km. Find the ratio of the average radius of the earth's orbit to the radius of the Hydrogen atom.

OpenStudy (paxpolaris):

http://www.lmgtfy.com/?q=1.5E8+km+%2F+(0.37+angstroms) \[\Large 4.1\times 10^{21} : 1\]

OpenStudy (paxpolaris):

\[\large {1.5\times 10^8km \over 0.37Å} ={1.5\times 10^{11}m \over 0.37\times 10^{-10}m}={4.1\times 10^{21} \over 1}\]